### v1

@@ 6.2 Experimental Results
Among the evaluated models, EgoGPT and Vinci are trained explicitly on egocentric datasets. Experimental results in Table 7 demonstrate that their classification performance on the CaptainCook4D dataset approaches random guessing, indicating limited detection capability. Other models, except Just Ask, exhibit only marginally better results. In terms of precision, these assistants significantly underperform compared to specialized error detection models like AMNAR, reflecting a limited understanding of error identification and limiting their suitability for direct application in error detection tasks. Further analysis reveals a consistent bias toward classifying observed segments as correct, with prompt adjustments yielding slight improvement. This deficiency likely stems from insufficient exposure to error-containing data during training, impairing the models’ understanding and detection of procedural mistakes. Moreover, unlike standard anomaly detection, procedural error detection requires comprehensive video understanding, as contextually correct actions may still constitute errors.
@@
Experimental results on EgoOops and EgoPER show that only EgoGPT, which was trained on egocentric datasets, effectively understands errors, while other models yield low precision scores. Vinci demonstrates some error detection for egocentric videos. However, under identical token length and prompt settings, it tends to generate lengthy analytical responses. This hampers the accurate calculation of confidence scores. For the EDA metric, some approaches surpass 70, but this does not imply stronger classification ability than EgoPED and AMNAR. Instead, models generally classify the majority of segments as correct, where performance relies heavily on the dataset’s ratio of positive samples. Thus, a higher proportion of correct segments yields higher EDA scores, but this does not mean these models outperform EgoPED or AMNAR in classification ability.
@@
In summary, through carefully designed experiments evaluating recent visual question answering models, we conclude that current VQA models cannot be directly applied to egocentric procedural error detection tasks. Their understanding of errors is limited, indicating their overall video comprehension requires further improvement.
@@
In the experimental evaluation of egocentric procedural learning shown in Table 8, the F1 and IoU metrics are utilized to assess the model’s capability. The F1 score evaluates step recognition accuracy, while the IoU quantifies the temporal overlap between predicted steps and ground-truth segments.
@@
Analysis of both PC assembly and disassembly tasks indicates that providing models with procedural knowledge does not yield satisfactory step recognition performance. However, in the context of PC disassembly and known procedural steps, Video-LLaMA2 outperforms two existing methods in step recognition accuracy, but still falls short of the state-of-the-art.
@@
Regarding the IoU metric, LLaVA-OneVision demonstrates improved temporal localization accuracy in the PC assembly task, surpassing one method but still requiring enhancement. Video-LLaMA2 also attains strong IoU results in PC disassembly. Although EgoGPT and Vinci are pre-trained on egocentric datasets, they do not show marked advantages. As they lack task-specific fine-tuning for egocentric procedural error detection or learning, definitive comparisons with other models are unwarranted. Further rigorous experiments are needed to assess their capabilities thoroughly.
@@
Table 8: Experimental results in egocentric procedural learning, the part highlighted in purple is the result of the model we selected, bold text represents the best performance on this metric in each part.
Method Years F1 IoU Task in EgoproceL F1 IoU Task in EgoproceL
CnC [bansal2022my] ECCV 2022 25.1 12.8 PC Assembly 27.0 14.8 PC Disassembly
OPEL [chowdhury2024opel] NeurIPS 2024 33.7 17.9 32.2 16.9
RGWOT [mahmood2025procedure] arXiv 2025 43.6 28.0 45.9 30.1
Vinci [huang2024vinci] arXiv 2024 14.1 7.5 27.2 14.7
Video-LLaVA [lin2023video] arXiv 2024 22.9 12.4 10.0 4.9
Video-LLaMA2 [cheng2024videollama] arXiv 2024 12.0 6.9 35.8 21.6
LLaVA-OneVision [li2024llava] arXiv 2024 22.2 13.0 16.8 8.8
EgoGPT [yang2025egolife] CVPR 2025 6.6 3.7 8.9 4.3
@@ 7 Challenges And Outlook
Our extensive experiments reveal that current vision-language models, including egocentric AI assistants, are insufficient for supporting procedural tasks. This highlights the necessity for further innovation to enable effective AI facilitation of such processes. Although notable progress has been made, egocentric vision tasks face ongoing challenges as they transition from controlled settings to real-world applications.
Data scarcity and bias. It impedes development in this domain. Existing egocentric video datasets are scarce, lack diversity, and provide limited annotation detail, constraining both error detection and procedural learning. While many procedural learning datasets are captured from third-person perspectives, they are ill-suited for egocentric tasks. Most available egocentric error detection datasets also cover narrow ranges of activities, failing to support comprehensive AI training for a wide range of real-world tasks. Additionally, cultural and individual biases present in current datasets further hinder robust generalization.
Limited understanding of long-term procedural tasks. Current models are struggling to capture logical and temporal dependencies in procedural videos, failing to distinguish true procedural errors or provide actionable explanations. They encounter difficulties in computational efficiency, semantic understanding, and the effective integration of multimodal signals, which hampers comprehensive video interpretation. Existing approaches, such as error detection, rely on a foundational model comprehension of procedural steps. Future research should prioritize adapting advanced long-video understanding frameworks to egocentric contexts and enhancing multimodal fusion for deeper procedural task analysis and improved model support.
Significantly depend on manual annotations and lack real-time capability. For AI assistants to effectively support daily life, real-time operation is essential. Human annotation dependence limits the practicality of real-world applications, constrains real-time detection, and narrows application domains. Currently, few approaches perform effectively in real-time online contexts. Providing AI assistants with instantaneous assistance and rapid responses remains a critical challenge. The methodology of predicting subsequent steps from previous actions or eye-tracking data requires further scholarly exploration [mazzamuto2025gazing][flaborea2024prego]. Integrating this predictive framework with the reasoning and learning abilities of LLMs may improve real-time assistance performance.
@@
Egocentric procedural planning [liu2023egocentric][chen2025planning]. It focuses on analyzing completed actions to generate context-appropriate subsequent steps. This task requires advanced comprehension and reasoning over egocentric video streams, enabling models to autonomously select optimal next actions based on real-time context.
Egocentric reasoning [vinod2025egovlm][lee2025towards][peirone2025hier]. It is an emerging area aiming for deep causal, temporal, and logical inference over egocentric video content. Recent research prioritizes understanding action relationships, accurate episodic memory retrieval, social interaction, and hand-object analysis [su2024care][su2025annexe][deng2025egocentric], moving beyond simple associative learning.
Integrating techniques from these domains can significantly enhance models’ comprehension of procedural tasks, leading to comprehensive performance improvements and enabling AI assistants to better support our daily lives.
@@ 8 Conclusion
This work conceptualizes an egocentric procedural AI assistant (EgoProceAssist), structuring the field into three core areas: egocentric procedural error detection, egocentric procedural learning, and egocentric procedural question answering. Existing literature is systematically reviewed and categorized. A novel taxonomy is introduced, providing a systematic classification and synthesis of existing methodological approaches. Subsequently, a thorough examination of prevalent datasets and evaluation metrics used in assessing the three principal tasks is presented. In addition, a new experimental paradigm is devised to rigorously evaluate the performance of generative AI assistants in the domains of egocentric procedural error detection and procedural learning. Empirical findings substantiate that extant AI assistants are presently insufficient for the direct support of procedural tasks. Lastly, the study delineates the persistent challenges and inherent limitations in egocentric vision research, with a particular emphasis on the construction of intelligent agents, and elucidates prospective avenues that may significantly influence subsequent advances in the field. We hope this work informs and inspires further research within the community.
@@ 9 Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.
@@ 10 Data Availability Statement
The datasets used for the supplementary experiment are all publicly available; however, to obtain EgoPER, you must email the author. The datasets we show in Table 2,  4 and  6 are all publicly available. The experiment data shown in Table 1,  3 and  5 are all publicly available. To make it easier to use, we have summarized them in https://github.com/z1oong/Building-Egocentric-Procedural-AI-Assistant.

### v2

@@ 7.2 Experimental Results
Among the evaluated models, EgoGPT and Vinci are trained explicitly on egocentric datasets. Experimental results in Table 10 demonstrate that their classification performance on the CaptainCook4D dataset approaches random guessing, indicating limited detection capability. Other models, except Just Ask, exhibit only marginally better results. In terms of precision, these assistants significantly underperform specialized error detection models like AMNAR, reflecting a limited understanding of error identification and limiting their suitability for direct application in error detection tasks. Further analysis reveals a consistent bias toward classifying observed segments as correct, with prompt adjustments yielding slight improvement. This deficiency likely stems from insufficient exposure to error-containing data during training, impairing the models’ understanding and detection of procedural mistakes. Moreover, unlike standard anomaly detection, procedural error detection requires a comprehensive understanding of video content, as contextually correct actions may still constitute errors.
@@
Experimental results on EgoOops and EgoPER show that only EgoGPT, trained on egocentric datasets, effectively understands errors, whereas other models yield low precision scores. Multimodal LLMs with strong reasoning capabilities, such as Ming-Lite-Omni, achieve high AUC, F1, and Recall scores, indicating strong overall detection performance. However, under identical token length and prompt settings, it tends to generate lengthy analytical responses. This hampers the accurate calculation of confidence scores. For the EDA metric, some approaches surpass 70, but this does not imply stronger classification ability than EgoPED and AMNAR. Instead, models generally classify the majority of segments as correct, where performance relies heavily on the dataset’s ratio of positive samples. Thus, a higher proportion of correct segments yields higher EDA scores, but this does not mean these models outperform EgoPED or AMNAR in classification ability. Under the in-context learning strategy, added context and examples in the prompt play a clear role, providing task-specific structure, clear error definitions, and simple examples of each error type, leading to substantial improvements in multiple metrics for the selected models. This demonstrates that prompt-based strategies can enhance models’ capability for procedural error detection to some extent, although there remains a performance gap compared with EgoPED and AMNAR, which in turn indicates a direction for future improvement.
@@
In summary, through carefully designed experiments evaluating recent VQA models, we conclude that current models cannot be directly applied to egocentric procedural error detection tasks. Their understanding of errors is limited, indicating their overall video comprehension requires further improvement.
@@
In the experimental evaluation of egocentric procedural learning shown in Table 11, F1 and IoU metrics are utilized to assess the model’s capability. F1 score evaluates step recognition accuracy, while the IoU quantifies the temporal overlap between predicted steps and ground-truth segments.
@@
Analysis of both PC assembly and disassembly tasks indicates that providing models with procedural knowledge does not yield satisfactory step recognition performance. However, in the context of PC disassembly and known procedural steps, Video-LLaMA2 outperforms two existing methods in step recognition accuracy, but still falls short of the state-of-the-art methods.
@@
Regarding the IoU metric, LLaVA-OneVision demonstrates improved temporal localization accuracy in the PC assembly task, surpassing one method but still requiring enhancement. Video-LLaMA2 also attains strong IoU results in PC disassembly. Although EgoGPT and Vinci are pre-trained on egocentric datasets, they do not show marked advantages. As they lack task-specific fine-tuning for egocentric procedural error detection or learning, definitive comparisons with other models are unwarranted. Further rigorous experiments are needed to assess their capabilities thoroughly. Finally, in the error detection task, Video-LLaVA achieves the shortest average inference time, while in procedural learning, Vinci is the fastest. Still, both operate on the order of seconds, which may be insufficient for providing real-time and effective support.
@@
Table 11: Experimental results in egocentric procedural learning, the part highlighted in purple is the result of the model we selected, bold text represents the best performance on this metric in each part, and Time is the average inference time.
Method Years Time F1 IoU Task Time F1 IoU Task
CnC [bansal2022my] ECCV 2022 - 25.1 12.8 PC Assembly - 27.0 14.8 PC Disassembly
OPEL [chowdhury2024opel] NeurIPS 2024 - 33.7 17.9 - 32.2 16.9
RGWOT [mahmood2025procedure] arXiv 2025 - 43.6 28.0 - 45.9 30.1
Video-LLaVA [lin2024video] EMNLP 2024 5.34 22.9 12.4 6.03 10.0 4.9
Video-LLaMA2 [cheng2024videollama] arXiv 2024 4.85 12.0 6.9 5.44 35.8 21.6
Vinci [huang2025vinci] IMWUT 2025 4.75 14.1 7.5 4.83 27.2 14.7
LLaVA-OneVision [lillava] TMLR 2025 7.18 22.2 13.0 7.36 16.8 8.8
EgoGPT [yang2025egolife] CVPR 2025 6.86 6.6 3.7 7.57 8.9 4.3
Qwen2.5-Omni-7B [xu2025qwen2] arXiv 2025 15.0 9.6 5.6 15.8 1.5 0.8
Ming-Lite-Omni [ai2025ming] arXiv 2025 17.4 9.3 5.9 16.6 7.6 4.5
@@ 8 Challenges and Outlook
8.1 Challenges
Our extensive experiments reveal that current vision-language models, including egocentric AI assistants, are insufficient for supporting procedural tasks. This highlights the necessity for further innovation to enable effective AI facilitation of such processes. Although notable progress has been made, egocentric vision tasks face challenges as they transition from controlled settings to real-world applications.
System-level bottlenecks, latency, and deployment constraints. Although we have analyzed real-time and streaming video understanding to reduce latency, major challenges remain in deploying models on wearable devices [lane2016deepx]. Many VLMs are too large for current hardware and, even when deployable, still suffer from latency, overheating, and related issues. Larger models usually deliver better reasoning and performance, and added functionality increases resource demands. Balancing model size with deployment constraints is thus a central challenge for realizing practical, real-world AI assistants.
Data scarcity and bias. Existing egocentric video datasets are scarce, lack diversity, and provide limited annotation detail, constraining both error detection and procedural learning. While many procedural learning datasets are captured from third-person perspectives, they are ill-suited for egocentric tasks. Most available egocentric error detection datasets also cover narrow ranges of activities, failing to support comprehensive AI training for a wide range of real-world tasks. Additionally, cultural and individual biases present in current datasets further hinder robust generalization. Human annotation dependence limits the practicality of real-world applications, constrains real-time detection, and narrows application domains. Recent egocentric generation methods, such as diffusion-based models like LEGO [lai2024lego], can be used to augment datasets and mitigate data scarcity, annotation gaps, and inherent biases.
Limited understanding of long-term procedural tasks. Current models are struggling to capture logical and temporal dependencies in procedural videos, failing to distinguish true procedural errors or provide actionable explanations. They encounter difficulties in computational efficiency, semantic understanding, and the effective integration of multimodal signals, which hampers comprehensive video interpretation. Existing approaches, such as error detection, rely on a foundational model comprehension of procedural steps. Future research should prioritize adapting advanced long-video understanding frameworks [goletto2024amego, plizzari2025spatial] to egocentric contexts and enhancing multimodal fusion [xu2025qwen2, ai2025ming, xu2025qwen3omnitechnicalreport, xie2024miniomni2opensourcegpt4ovision, meituanlongcatteam2025longcatflashomnitechnicalreport] for deeper procedural task analysis and improved model support.
8.2 Egocentric Procedural Reasoning, Planning, and Action Anticipation
@@
Egocentric procedural reasoning [vinod2025egovlm, lee2025towards, peirone2025hier]. It is an emerging area aiming for deep causal, temporal, and logical inference over egocentric video content. Recent research prioritizes understanding action relationships, accurate episodic memory retrieval, social interaction, and hand-object analysis [su2024care, su2025annexe, deng2025egocentric, wang2019object], moving beyond simple associative learning. Given the preceding results, current models and assistants show limited competence in procedural tasks. Improving their reasoning through in-context learning, chain-of-thought prompting, or multi-agent systems with deliberate reasoning and tool-use capabilities may effectively address these procedural challenges.
Egocentric procedural planning and action anticipation. Recent work [zhao2023antgpt, fang2024egocentric, kulkarni2025egovita, liu2023egocentric, rhinehart2017first, rodin2024action] increasingly strengthens models’ procedural understanding by recognizing past and ongoing actions to infer and predict future actions and action sequences. For EgoProceAssist, this paradigm is not only a key foundation but also a promising route to tackling procedure-centric tasks. Prior studies such as PREGO [flaborea2024prego], TI-PREGO [plini2025ti], and AMNAR [huang2025modeling] detect errors by predicting the next plausible action at each timestep, while GC [mazzamuto2025gazing] predicts future gaze from eye-movement trajectories and flags errors via gaze deviations, achieving accurate error detection with real-time responsiveness without processing the full video. These results show that anticipation provides a novel mechanism for error detection and enables inferring user intent from signals such as gaze, which is crucial for building practical assistants. In procedural learning and QA, anticipation supports online updating of step sequences, offering new ways to monitor task progress and verify results. By anticipating user intent, the assistant can proactively deliver support, plan feasible action paths, and provide targeted guidance. Consequently, egocentric procedural planning and action anticipation emerge as a central direction that warrants sustained and systematic investigation.
8.3 Edge AI, Model Distillation, and Cloud-Edge Architectures
Recent studies employ lightweight backbone networks with feature compression and streaming video processing to support deployment on smart glasses and to exploit fog and cloud services. Future work can prioritize edge AI and on-device inference. Lightweight architectures and hardware-aware optimizations are key to deploying AI assistants on resource-constrained devices. Model compression techniques such as pruning [zhu2017prune], quantization [han2015deep], and knowledge distillation [hong2022analysis], have been applied to VLMs to reduce size and latency with minimal accuracy loss, including distilling large VLMs into smaller, task-specific models. Cloud–edge collaborative architectures [yu2025cloud] are also promising. Hybrid systems run lightweight models on-device for latency-critical tasks (e.g., error detection) and offload complex reasoning to the cloud, balancing latency and computational load.
Integrating techniques from these domains can significantly enhance models’ comprehension of procedural tasks, leading to substantial performance improvements and enabling AI assistants to better support our daily lives.
@@ 9 Conclusion
This work conceptualizes an egocentric procedural AI assistant (EgoProceAssist), structuring the field into three core areas: egocentric procedural error detection, egocentric procedural learning, and egocentric procedural question answering, along with two enabling dimensions: real-time and streaming video understanding, and proactive interaction in procedural contexts. Existing literature is systematically reviewed and categorized. A novel taxonomy is introduced that provides a systematic classification and synthesis of existing methodological approaches. Subsequently, a thorough examination of the prevalent datasets and evaluation metrics used to assess the three principal tasks is presented. In addition, a new experimental paradigm is devised to rigorously evaluate the performance of generative AI assistants in the domains of egocentric procedural error detection and procedural learning. Empirical findings substantiate that extant AI assistants are presently insufficient for the direct support of procedural tasks. Lastly, the study delineates the persistent challenges and inherent limitations in egocentric vision research, with a particular emphasis on the construction of intelligent agents, and elucidates prospective avenues that may significantly influence subsequent advances in the field. We hope this work informs and inspires further research within the community.
@@ 10 Acknowledgment
The research work described in this paper was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust. This research received partially support from the Global STEM Professorship Scheme from the Hong Kong Special Administrative Region.
@@ 11 Data Availability Statement
The datasets used for the supplementary experiment are all publicly available; however, to obtain EgoPER, you must email the author. The datasets we show in Table 2,  4,  6,  8, and  9 are all publicly available. The experiment data shown in Table 1,  3,  5, and  7 are all publicly available. To make it easier to use, we have summarized them in https://github.com/z1oong/Building-Egocentric-Procedural-AI-Assistant.